# Learning Useful Representations of Recurrent Neural Network Weight Matrices

### v1

@@ 6.2 Downstream Property Prediction
For comparison, we also trained models for property prediction in a purely supervised manner. The setup is the same as for the pre-trained scenario, consisting of an RNN weight encoder followed by an MLP predictor. However, in this case, the encoder is not pre-trained but is randomly initialized and trained end-to-end with the predictor. Our primary interest lies in the generalization capabilities of the pre-trained $E_{\phi}$, rather than the predictor itself. Consequently, the training data for the predictor in both supervised and pre-trained settings includes half of the OOD test set. We place particular emphasis on task prediction (specifically, the type of formal language or the rotation of the MNIST dataset). This is because it demands a genuine understanding of the RNN’s function, which goes beyond mere high-level statistical analysis of the weights. Figure 8 (top) demonstrates that for task prediction in the formal languages dataset (OOD split), the pre-trained representations from an interactive probing encoder significantly outperform those from other encoders, as well as purely supervised models. For the sequential MNIST dataset (Figure 8, bottom), both types of probing encoders surpass other architectures. In the supervised scenario, the non-interactive probing encoder excels, but the interactive version does not perform as well. We hypothesize that this is due to the limited information provided by supervised training compared to self-supervised pre-training, which does not offer sufficient feedback for the interactive encoder to develop effective probing sequences. Complete results for downstream predictions, both supervised and pre-trained, on validation and OOD test data are shown in Figures 18 and 19 in the Appendix. Although these results do not definitively favor any single encoder architecture, they can be summarized as follows: for the Formal Languages dataset, on both task and performance prediction, interactive probing generally yields the best results, particularly in the pre-trained setting. For the Sequential MNIST dataset, non-interactive probing performs best in the supervised setting, while both probing architectures are effective when pre-trained. In all other property predictions (accuracy, generalization gap, training step), DWSNet performs most consistently. For a more detailed discussion of the results, please see Appendix A.
@@ 7 Conclusion and Future Work
This work establishes a foundation for numerous future applications. For instance, the techniques introduced here can be applied in the context of reinforcement learning within partially observable environments, where they can facilitate policy representation and improvement (see e.g., Raileanu et al. (2020); Faccio et al. (2020, 2022b)), as well as exploration and skill discovery (Herrmann et al., 2022). Additionally, RNN weight representations could be beneficial in meta-learning and few-shot learning scenarios. Our functionalist probing approaches have the potential to extract information not only from small model populations but also from large foundational models. This can be achieved by conditioning either the encoder or the foundational model itself on a specific task, a concept that aligns with discussions in previous research (Schmidhuber, 2015; Zeng et al., 2022; Zhuge et al., 2023).
@@ Acknowledgements
Appendix A Additional Discussion
Third, an encoder architecture capable of handling the complex structure of RNN weights as input is crucial. Except for Layer-Wise statistics, mechanistic encoders can approximate any function within the weight space. Nonetheless, in practice, functionalist approaches outperform them both in self-supervised training objectives and task prediction. Our findings confirm the theoretical prediction that interactive probing is more efficient for certain datasets (Proposition 3.3), as evidenced by the results on the Formal Languages datasets. However, interactive probing may suffer from training stability issues in some instances. When the interactive property is unnecessary or too complex to train, such as with the sequential MNIST dataset, non-interactive probing may yield better results. Additionally, functionalist encoders offer the advantage of being agnostic to the precise architecture of $f_{\theta}$. In more straightforward tasks like supervised property prediction, mechanistic encoders like DWSNet excel.
Appendix B Theoretical results
@@
According to Lemma B.1, it is possible to split any set $G\subseteq D,|G|\geq 2$ into two nonempty, non-overlapping subsets: $G_{a}:=\{f\in G|f(x_{j})=f_{a}(x_{j})\}$ and $G_{b}:=\{f\in G|f(x_{j})\neq f_{a}(x_{j})\}$ for some $f_{a}\in G$ and $x_{j}\in\mathbb{N}$. Any resulting subset that has at least two members can be split again using the same procedure with a different probing input $x_{j+1}$. Starting from the full set $D$, it is possible to construct a binary tree (see Figure 9) where the leaves are subsets of $D$ containing exactly one uniquely identified function. The branching (i.e., non-leaf) nodes correspond to the splitting operation, which involves observing the output of a specific probing input $x_{j}$.
@@
Appendix C DWSNet Details
Figure 10: All connections from and to a single weight of an RNN in a input to hidden weight matrix (left) or hidden to hidden weight matrix (right). They must all be taken into account to ensure equivariance of DWSNets for RNN weights.
@@ Appendix C DWSNet Details
Appendix D Dataset Details
The process for creating the Formal Languages and Sequential MNIST RNN weight datasets is consistent across all instances. In each of the 1000 training iterations, a new LSTM network is initialized, and a task is chosen at random. This model undergoes training for a total of 20000 steps using the AdamW optimizer (Loshchilov & Hutter, 2017), with a weight decay of $10^{-4}$ and a batch size of 32. We implement a piece-wise linear learning rate schedule that adjusts the learning rate between $[0.01,0.003,0.003]$ at steps $[0,10000,20000]$. The model weights are saved as a datapoint at the training steps $[0,100,200,500,1000,2000,5000,10000,20000]$, along with 100 rollout sequences and various attributes (such as the task, training step, and performance metrics, including Sequential MNIST validation and training loss).
Appendix E Experiment Details
@@
Table 10: Interactive Probing Hyperparameters
Appendix F Additional Results
F.1 Probing Sequence Number and Length
@@
Figure 14: Performance of $f_{\theta}$’s original performance on sequential MNIST classification vs. the performance of $A_{\xi}$’s emulation based on $E_{\phi}(\theta)$.
@@
The probing encoders presented in this work can utilize a fixed number of parallel probing sequences. If not specified otherwise, 8 probing sequences are used. Figures 12 and 12 demonstrate that using more probing sequences correlates with lower loss. Interactive probing proves highly effective for RNNs applied to the Formal Languages dataset, though it faces challenges with training stability. In principle, one probing sequence is sufficient to solve the task, as evidenced by the top-performing model among 15 seeds achieving very low loss regardless of the number of probing sequences. Nonetheless, training stability improves with an increase in probing sequences.
Figure 15 displays the probing sequences generated by different Interactive Probing encoders for a specific RNN that produces strings from languages $L_{-3,-1,1}$ the encoders for this figure learn only one probing sequence). A length of $7$ is notably insufficient, failing to learn any meaningful probing sequence. However, at sequence lengths of $12$, $22$, and $42$, the encoder successfully learns to probe actual strings of varying lengths from $L_{-3,-1,1}$. Note that while there is a string of length less than $12$ for this particular language, this does not apply to all languages in the dataset.
F.2 Hidden Neuron Permutation Invariance
@@
Figures 17 and 17 illustrate the invariance of different encoder architectures towards correct (first column) and incorrect (other columns) hidden neuron permutations. Ideally, invariance should be present only in the first column. However, the Layerwise Statistics encoder shows invariance to both correct and incorrect permutations. In contrast, the Flattened Weights and Parameter Transformer encoders exhibit no invariance, neither before (Figure 17) nor after training (Figure 17). During training, DWSNet becomes less invariant to wrong and incomplete permutations. The methodology of the plot is explained below.
@@
F.3 Downstream Performance
Table 11 shows the Pearson correlation coefficients between pre-training validation loss, as defined in Equation 1, and the downstream prediction losses for different properties, across all encoder architectures and random seeds. This analysis offers insight into the alignment between the pre-training objective and downstream applications. There is a high correlation with formal language task prediction and with sequential MNIST performance prediction as well as in-distribution task prediction. A lower correlation does not necessarily imply that pre-training is less effective; it may simply indicate that moderate emulation performance is sufficient for good downstream prediction results. Figures 18 and 19 provide the complete results of downstream predictions for Formal Languages and Sequential MNIST data.
@@ F.3 Downstream Performance
Figure 19: Sequential MNIST downstream performance on task, performance, generalization gap and training step prediction.
F.4 Learned Embedding Spaces
@@
Figure 21: Visualization of the embedding spaces for formal languages (colored by task and return).
Figure 22: Visualization of the embedding spaces for formal languages (colored by training step and generalization gap).

### v2

@@ 6.2 Downstream Property Prediction
For comparison, we also trained models for property prediction in a purely supervised manner. The setup is the same as for the pre-trained scenario, consisting of an RNN weight encoder followed by an MLP predictor. However, in this case, the encoder is not pre-trained but is randomly initialized and trained end-to-end with the predictor. Our primary interest lies in the generalization capabilities of the pre-trained $E_{\phi}$, rather than the predictor itself. Consequently, the training data for the predictor in both supervised and pre-trained settings includes half of the OOD test set. We place particular emphasis on task prediction (specifically, the type of formal language or the rotation of the MNIST dataset). This is because it demands a genuine understanding of the RNN’s function, which goes beyond mere high-level statistical analysis of the weights. Figure 8 (top) demonstrates that for task prediction in the Formal Languages dataset (OOD split), the pre-trained representations from an interactive probing encoder significantly outperform those from other encoders, as well as purely supervised models. For the Sequential MNIST dataset (Figure 8, bottom), both types of probing encoders surpass other architectures. In the supervised scenario, the non-interactive probing encoder excels, but the interactive version does not perform as well. We hypothesize that this is due to the limited information provided by supervised training compared to self-supervised pre-training, which does not offer sufficient feedback for the interactive encoder to develop effective probing sequences. Complete results for downstream predictions, both supervised and pre-trained, on validation and OOD test data are shown in Figures 19 and 20 in the Appendix. Although these results do not definitively favor any single encoder architecture, they can be summarized as follows: for the Formal Languages dataset, on both task and performance prediction, interactive probing generally yields the best results, particularly in the pre-trained setting. For the Sequential MNIST dataset, non-interactive probing performs best in the supervised setting, while both probing architectures are effective when pre-trained. In all other property predictions (accuracy, generalization gap, training step), DWSNet performs most consistently.
@@
7 Discussion
Third, an encoder architecture capable of handling the complex structure of RNN weights as input is crucial. Except for Layer-Wise statistics, mechanistic encoders can approximate any function within the weight space. Nonetheless, in practice, functionalist approaches outperform them both in self-supervised training objectives and task prediction. Our findings confirm the theoretical prediction that interactive probing is more efficient for certain datasets (Proposition 3.3), as evidenced by the results on the Formal Languages datasets. However, interactive probing may suffer from training stability issues in some instances. This issue might be mitigated in the future through different training methods, loss functions, or regularizations. When the interactive property is unnecessary or too complex to train, such as with the Sequential MNIST dataset, non-interactive probing may yield better results. Additionally, functionalist encoders offer the advantage of being agnostic to the precise architecture of $f_{\theta}$. In more straightforward tasks like supervised property prediction, mechanistic encoders like or Layer-Wise Statics or DWSNet excel.
To date, our work is limited to relatively small RNNs. All approaches presented here can in principle be adapted in a straightforward way to other architectures, such as state space models (Gu et al., 2021), deep equilibrium models (Bai et al., 2019), or transformers (Vaswani et al., 2017). For scaling to larger models, we see the greatest potential in functionalist probing approaches. These are, in principle, agnostic to the architecture and size of the RNN (or any general sequence model), provided it is differentiable. By employing policy gradient methods, we could even process models that are impossible or prohibitively expensive to differentiate.
@@ 8 Conclusion and Future Work
This work establishes a foundation for numerous future applications. For instance, the techniques introduced here can be applied in the context of reinforcement learning within partially observable environments, where they can facilitate policy representation and improvement (see e.g., Raileanu et al. (2020); Faccio et al. (2020, 2022b)), as well as exploration and skill discovery (Herrmann et al., 2022). Additionally, RNN weight representations could be beneficial in meta-learning and few-shot learning scenarios. Another field where our work might be useful is the mechanistic interpretability of sequence models (e.g., (Weiss et al., 2018; Olsson et al., 2022)). Typically, it involves a lot of labor-intensive reverse engineering. Training models to create meaningful representations of RNN (or general sequence model) weights may assist or even partially automate these efforts. Our functionalist probing approaches have the potential to extract information not only from populations of small models but also from large foundational models. This can be achieved by conditioning either the encoder or the foundational model itself on a specific task, a concept that aligns with discussions in previous research (Schmidhuber, 2015; Zeng et al., 2022; Zhuge et al., 2023).
@@ Acknowledgements
Appendix A Theoretical results
@@
According to Lemma A.1, it is possible to split any set $G\subseteq D,|G|\geq 2$ into two nonempty, non-overlapping subsets: $G_{a}:=\{f\in G|f(x_{j})=f_{a}(x_{j})\}$ and $G_{b}:=\{f\in G|f(x_{j})\neq f_{a}(x_{j})\}$ for some $f_{a}\in G$ and $x_{j}\in\mathbb{N}$. Any resulting subset that has at least two members can be split again using the same procedure with a different probing input $x_{j+1}$. Starting from the full set $D$, it is possible to construct a binary tree (see Figure 9) where the leaves are subsets of $D$ containing exactly one uniquely identified function. The branching (i.e., non-leaf) nodes correspond to the splitting operation, which involves observing the output of a specific probing input $x_{j}$.
@@
Appendix B DWSNet Details
Figure 10: All connections from and to a single weight of an RNN in an input-to-hidden weight matrix (left) or hidden-to-hidden weight matrix (right). They must all be taken into account to ensure the equivariance of DWSNets for RNN weights.
@@ Appendix B DWSNet Details
Appendix C Dataset Details
The process for creating the Formal Languages and Sequential MNIST RNN weight datasets is consistent across all instances. In each of the 1000 training iterations, a new LSTM network is initialized, and a task is chosen at random. This model undergoes training for a total of 20000 steps using the AdamW optimizer (Loshchilov & Hutter, 2017), with a weight decay of $10^{-4}$ and a batch size of 32. We implement a piece-wise linear learning rate schedule that adjusts the learning rate between $[0.01,0.003,0.0003]$ at steps $[0,10000,20000]$. The model weights are saved as a datapoint at the training steps $[0,100,200,500,1000,2000,5000,10000,20000]$, along with 100 rollout sequences and various attributes (such as the task, training step, and performance metrics, including Sequential MNIST validation and training loss).
Appendix D Experiment Details
@@
Appendix E Additional Results
Figure 11 shows the emulation performance of different emulators $A_{\xi}$ on the Formal Languages datasets. We observe overfitting on the training data of the non-interactive probing and parameter transformer encoders. Note, however, that this occurs despite the use of early stopping, meaning the models evaluated are those with the lowest validation loss.
Figure 12 shows the emulation performance on the Sequential MNIST datasets. There we can observe an interesting effect: in the Parameter Transformer setup, $A_{\xi}$ always classifies the digits with high accuracy, even when conditioned on low-performing RNNs. We suggest the following explanation: from Figures 20, 22 and 23 it is clear that for Sequential MNIST, the parameter transformer fails to learn informative representations. What does the emulator $A_{\xi}$ learn in this case? For Sequential MNIST, the target distributions $y$ typically assign a high probability to the correct digit. Therefore, the emulator is incentivized to learn a rotation-invariant MNIST classifier that will perform well on any input, regardless of the conditioning. However, there is an additional nuance: the reverse KL divergence loss function prefers high entropy output distributions from the emulator. This means that while the emulator’s performance in terms of classification loss would be poor (since high entropy distributions generally result in high losses), its performance in terms of classification accuracy can still be good even with a high entropy distribution, as long as the correct digits have a probability that is at least slightly higher than the other ones.
Figure 11: Formal Languages original performance vs. emulated performance. Training (left), validation (middle), and OOD test data (right).
Figure 12: Sequential MNIST original performance vs. emulated performance. Training (left), validation (middle), and OOD test data (right)
E.1 Probing Sequence Number and Length
The probing encoders presented in this work can utilize a fixed number of parallel probing sequences. If not specified otherwise, 8 probing sequences are used. Figures 15 and 15 demonstrate that using more probing sequences correlates with lower loss. Interactive probing proves highly effective for RNNs applied to the Formal Languages dataset, though it faces challenges with training stability. In principle, one probing sequence is sufficient to solve the task, as evidenced by the top-performing model among 15 seeds achieving very low loss regardless of the number of probing sequences. Nonetheless, training stability improves with an increase in probing sequences.
Figure 16 displays the probing sequences generated by different Interactive Probing encoders for a specific RNN that produces strings from languages $L_{-3,-1,1}$ (the encoders for this figure learn only one probing sequence). A length of $7$ is notably insufficient, failing to learn any meaningful probing sequence. However, at sequence lengths of $12$, $22$, and $42$, the encoder successfully learns to probe actual strings of varying lengths from $L_{-3,-1,1}$. Note that while there is a string of length less than $12$ for this particular language, this does not apply to all languages in the dataset.
@@ E.1 Probing Sequence Number and Length
E.2 Hidden Neuron Permutation Invariance
@@
Figures 18 and 18 illustrate the invariance of different encoder architectures towards correct (first column) and incorrect (other columns) hidden neuron permutations. Ideally, invariance should be present only in the first column. However, the Layerwise Statistics encoder shows invariance to both correct and incorrect permutations. In contrast, the Flattened Weights and Parameter Transformer encoders exhibit no invariance, neither before (Figure 18) nor after training (Figure 18). During training, DWSNet becomes less invariant to wrong and incomplete permutations. The methodology of the plot is explained below.
@@
E.3 Downstream Performance
Table 11 shows the Pearson correlation coefficients between pre-training validation loss, as defined in Equation 1, and the downstream prediction losses for different properties, across all encoder architectures and random seeds. This analysis offers insight into the alignment between the pre-training objective and downstream applications. There is a high correlation with formal language task prediction and with Sequential MNIST performance prediction as well as in-distribution task prediction. A lower correlation does not necessarily imply that pre-training is less effective; it may simply indicate that moderate emulation performance is sufficient for good downstream prediction results. Figures 19 and 20 provide the complete results of downstream predictions for Formal Languages and Sequential MNIST data.
@@ E.3 Downstream Performance
Figure 20: Sequential MNIST downstream performance on task prediction, performance prediction, generalization gap, and training step prediction.
E.4 Learned Embedding Spaces
@@
Figure 22: Visualization of the embedding spaces for Sequential MNIST (colored by task and return).
Figure 23: Visualization of the embedding spaces for Sequential MNIST (colored by training step and generalization gap).